# Strategic Analysis of Fair Rank-Minimizing Mechanisms with Agent Refusal Option

### v2

@@ 2 Model
We consider the model discussed by Kojima and Manea (2010) and Featherstone (2020). Let $\mathcal{A}$ and $\mathcal{O}$ be finite sets of agents and object types respectively. Each agent must be allocated one copy of an object whose type is an element of a set denoted by $\mathcal{O}$. Let $q_{o}\geq 1$ be the capacity of an object type $o\in\mathcal{O}$. We assume that there is a null object type denoted by $\varnothing\in\mathcal{O}$ satisfying $q_{\varnothing}\geq\left|\mathcal{A}\right|$. That is, the null object type can be assigned to all agents simultaneously. The null object type $\varnothing\in\mathcal{O}$ can be interpreted as the (best) outside option of an agent. In Section 4, we consider the situation where the null object type has the other characteristic. To avoid trivial cases, we assume $q_{o}<\left|\mathcal{A}\right|$ for all $o\in\mathcal{O\setminus}\left\{\varnothing\right\}$.
Let $\mathcal{P}$ be the set of all possible strict rankings (linear orders) of the elements in $\mathcal{O}$. Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th best one under $p_{a}$. For notational convenience, we let $R:\mathcal{O\times P}\rightarrow\left\{1,\cdots,\left|\mathcal{O}\right|\right\}$ be a rank function; that is, $p_{a}^{k}=o$ if and only if $R\left(o,p_{a}\right)=k$. Moreover, if $R\left(o,p_{a}\right)<R\left(\varnothing,p_{a}\right)$ (resp. $R\left(o,p_{a}\right)>R\left(\varnothing,p_{a}\right)$), then $o$ is said to be acceptable (resp. unacceptable) for $p_{a}$. Thus, note that $R\left(\varnothing,p_{a}\right)-1$ is the number of acceptable object types for $p_{a}$. Let $r_{a}\in\mathcal{P}$ be the true preference order of agent $a$. Especially, if $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, then $o$ is said to be truly acceptable for $a$.
@@
$x_{ao}\in\left[0,1\right]\text{, }\mathop{\displaystyle\sum}\limits_{o^{\prime}\in\mathcal{O}}x_{ao^{\prime}}=1\text{,\ and }\mathop{\displaystyle\sum}\limits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}\leq q_{o},$
@@
$\mathop{\displaystyle\sum}\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}\geq\mathop{\displaystyle\sum}\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}^{\prime}$ (1)
for all $k=1,2,\cdots,\left|\mathcal{O}\right|$, then agent $a$ with $p_{a}$ is said to weakly prefer (in the sense of first-order stochastic dominance) $x$ to $x^{\prime}$. Moreover, $a$ is said to strictly prefers $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|$.
@@
Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\mathop{\textstyle\sum}\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$. Any wasteful assignment for $p$ must not be ordinally efficient. [^1] Moreover, a wasteful assignment for $p$ must not be efficient in a weaker efficiency concept such as ex-post efficiency, which is discussed several previous studies such as Bogomolnaia and Moulin (2001). Thus, not being wasteful is a very weak requirement for efficiency.
@@
$RV\left(x,p\right)=\mathop{\displaystyle\sum}\limits_{a\in\mathcal{A}}\mathop{\displaystyle\sum}\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times x_{ao},$
@@
$x^{\ast}=\mathop{\displaystyle\sum}\limits_{i=1}^{I}\sigma^{i}y^{i}\in\mathcal{X}\text{ where }\mathop{\displaystyle\sum}\limits_{i=1}^{I}\sigma^{i}=1$
@@
Nikzad (2022) and Troyan (2024) consider a deterministic rank efficient assignment; that is, they focus only on a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$. However, by Corollary 1, such $y^{\ast}$ must also be rank efficient in our definition. Moreover, this result allows us to show that the mechanism introduced later is well-defined.
Finally, we introduce another technical result that is used to show our main results introduced later.
@@
Let $x\in\mathcal{X}$ be an assignment that is not wasteful for $p$ and $k$ be the smallest integer satisfying
$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}>\left|\mathcal{A}\right|.$ (2)
Then, $x_{ao}=0$ for all $o$ such that $R(o,p_{a})\geq k$.
The condition (2) means that there are sufficient number of copies of object types that ranked higher than $k$ in the preference ranking of agent $a$. Then, if the assignment is not wasteful, then $a$ must be assigned to an object type that ranked higher than $k$. Typically, in that case, the object types that ranked lower than $\varnothing$ in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
@@ 3 Rank-minimizing Mechanism
Second, we introduce a fairness property. First, we introduce a well-known one. A mechanism $f$ is said to satisfy weak equal treatment of equals (hereafter weak ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}=p_{a^{\prime}}$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$. This requires that two agents revealing the same preference must have the same assignment in the ex-ante sense. This property is called just “equal treatment of equals” by several previous studies such as Bogomolnaia and Moulin (2001) and Troyan (2024). In this study, we mainly use a slightly stronger one. A mechanism $f$ is said to satisfy equal treatment of equals (hereafter ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $11,\cdots,k-$ 1 implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$, where $k$ is the smallest integer satisfying
$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}\geq\left|\mathcal{A}\right|\text{.}$
By Lemma 2, as long as the assignment is not wasteful, any type that is ranked $k$th or lower in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than $k$th. Thus, if two individuals reveal rankings that are identical for positions above the $k$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
@@
Let $\mathcal{Y}^{\ast}\left(p\right)\subseteq\mathcal{Y}$ be the set of deterministic rank efficient assignments; that is, $y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right),$ $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$. Note that by Lemma 1, $\mathcal{Y}^{\ast}\left(p\right)\neq\emptyset$. Let
$f^{U}\left(p\right)=\mathop{\displaystyle\sum}\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}\text{ for all }p\in\mathcal{P}^{\left|\mathcal{A}\right|}$
@@
$RV\left(f^{U}\left(p\right),p\right)=\mathop{\displaystyle\sum}\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{1}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}RV\left(y^{\ast},p\right).$
Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank efficient for $p$. Trivially, $f^{U}$ satisfies ETE and thus is a fair rank-minimizing mechanism. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the uniform rank-minimizing mechanism, this is one of the most common rank-minimizing mechanisms at least in the literature.
@@
We discuss strategic problems of fair rank-minimizing mechanisms. We call the strategy of revealing their true preference truth-telling. Featherstone (2020) shows that any rank-minimizing mechanism does not satisfy strategy-proofness. That is, for any rank-minimizing mechanism $f$, there are $r_{a}\in\mathcal{P}$ and $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ such that $a$ with $r_{a}$ prefers $f\left(p\right)$ to $f\left(r_{a},p_{-a}\right)$. Therefore, in this paper, we consider a weaker strategic property.
We discuss whether truth-telling ($r_{a}$) can be strategically dominated by an untrue preference $p_{a}\neq r_{a}$. Let $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$. By Lemma 2, if (2) is satisfied for that $k$, then $p_{a}$ and $r_{a}$ are essentially equivalent under a rank-minimizing mechanism. Thus, we focus only on the case where (2) is not satisfied for $k$ defined above; that is,
$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}\left(=\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},r_{a})<k}q_{\bar{o}}\right)<\left|\mathcal{A}\right|$ (3)
is satisfied for that $k$. In other word, we consider the case where the number of object types that are more preferable than the $k$th best one in $p_{a}$ is insufficient to distribute to all agents. In this case, for some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and some fair rank-minimizing mechanism, $a$ is assigned to an object type that is less preferable or equal to the $k$th best one.
@@
Suppose that $f$ is a fair rank-minimizing mechanism. Let $p_{a}\neq r_{a}$, and $k$ be the smallest integer such that $p_{a}^{k}\neq r_{a}^{k}$. If (3) is satisfied for that $k$, then there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that agent $a$ (with $r_{a}$) strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
Next, we have the following result.
@@
This result implies that any fair rank-minimizing mechanism satisfies a weak strategic characteristic; that is, for each untrue preference of an agent $a$, there must exist a profile of preferences of agents other than $a$, such that $a$ is worse off by revealing it instead of $r_{a}$. Thus, if a fair rank-minimizing mechanism is used, any strategic manipulation by an agent carries some level of risk. This characteristic is quite weak, but in the subsequent section, we show that this is not satisfied when the agents can refuse their assignment and instead obtain the null object type.
@@
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$. Let $\psi=\left(o_{1},o_{2},\varnothing,o_{3}\right)$ and $\psi^{\prime}=\left(o_{1},o_{2},o_{3},\varnothing\right)$. A rank-minimizing mechanism $f$ is such that
@@
$f\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ $=$ $\mathop{\displaystyle\sum}\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$
if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\psi^{\prime},\psi,\psi\right)$ nor $\left(\psi,\psi^{\prime},\psi\right)$ nor $\left(\psi,\psi,\psi^{\prime}\right)$. That is, if the best type and the second best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third best one obtains $o_{1}$ with probability 00. Trivially, $f$ satisfies weak ETE. However, whether $f$ satisfies ETE or not is dependent on $q_{o_{1}}$ and $q_{o_{2}}$.
Suppose $r_{a_{1}}=\psi^{\prime}$ and this rank-minimizing mechanism is used. First, we assume $q_{o_{1}}=q_{o_{2}}=1$. In this case, $r_{a_{1}}=\psi^{\prime}$ is not strategically dominated by $\psi$ and any other preference orders, because
@@
On the other hand, suppose $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$. In this case, $a_{1}$ must have either $o_{1}$ or $o_{2}$ if it chooses either $\psi$ or $\psi^{\prime}$ and thus, the unique interest of $a_{1}$ is increasing the probability to obtain $o_{1}$. Thus, in this case, $r_{a_{1}}=\psi^{\prime}$ is strategically dominated by $\psi$. Therefore, even if a fair rank-minimizing mechanism is used, truth-telling may be strategically dominated by some other strategy. Moreover, in this case, $f$ does not satisfy ETE. First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing$,∅ , as long as the assignment is not wasteful. Thus, $\psi$ and $\psi^{\prime}$ are essentially the same in this case. However, $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}=0,$ but $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{2}}>0$ or $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{3}}>0$ must be satisfied. Therefore, in this case, $f$ does not satisfy ETE.

### v3

@@ 2 Model
We consider the model discussed by Kojima and Manea (2010) and Featherstone (2020). Let $\mathcal{A}$ and $\mathcal{O}$ be finite sets of agents and object types respectively. We assume $\left|\mathcal{A}\right|\geq 2$ and $\left|\mathcal{O}\right|\geq 2$. Each agent must be allocated one copy of an object whose type is an element of a set denoted by $\mathcal{O}$. Let $q_{o}\geq 1$ be the capacity of an object type $o\in\mathcal{O}$. We assume that there is a null object type denoted by $\varnothing\in\mathcal{O}$ satisfying $q_{\varnothing}\geq\left|\mathcal{A}\right|$. That is, the null object type can be assigned to all agents simultaneously. The null object type $\varnothing\in\mathcal{O}$ can be interpreted as the (best) outside option of an agent. In Section 4, we consider the situation where the null object type has the other characteristic. To avoid trivial cases, we assume $q_{o}\in\left[1,\left|\mathcal{A}\right|\right)$ for all $o\in\mathcal{O\setminus}\left\{\varnothing\right\}$.
Let $\mathcal{P}$ be the set of all possible strict rankings (linear orders) of the elements in $\mathcal{O}$. Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}^{\left|\mathcal{A}\right|}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th best one under $p_{a}$. For notational convenience, we let $R:\mathcal{O\times P}\rightarrow\left\{1,\cdots,\left|\mathcal{O}\right|\right\}$ be a rank function; that is, $p_{a}^{k}=o$ if and only if $R\left(o,p_{a}\right)=k$. Moreover, if $R\left(o,p_{a}\right)<R\left(\varnothing,p_{a}\right)$ (resp. $R\left(o,p_{a}\right)>R\left(\varnothing,p_{a}\right)$), then $o$ is said to be acceptable (resp. unacceptable) for $p_{a}$. Thus, $R\left(\varnothing,p_{a}\right)-1$ represents the number of acceptable object types for $p_{a}$. Let $r_{a}\in\mathcal{P}$ be the true preference order of agent $a$. Especially, if $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, then $o$ is said to be truly acceptable for $a$.
@@
$x_{ao}\in\left[0,1\right]\text{, }\sum\limits_{o^{\prime}\in\mathcal{O}}x_{ao^{\prime}}=1\text{,\ and }\sum\limits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}\leq q_{o},$
@@
$\sum\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}\geq\sum\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}^{\prime}$ (1)
for all $k=1,2,\cdots,\left|\mathcal{O}\right|$, then agent $a$ with $p_{a}$ is said to weakly prefer (in the sense of first-order stochastic dominance) $x$ to $x^{\prime}$. Moreover, $a$ is said to strictly prefers $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|-1$.
@@
Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\sum\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$. Any wasteful assignment for $p$ must not be ordinally efficient. [^1] Moreover, a wasteful assignment for $p$ must not be efficient in a weaker efficiency concept such as ex-post efficiency, which is discussed several previous studies such as Bogomolnaia and Moulin (2001). Thus, not being wasteful is a very weak requirement for efficiency.
@@
$RV\left(x,p\right)=\sum\limits_{a\in\mathcal{A}}\sum\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times x_{ao},$
@@
$x^{\ast}=\sum\limits_{i=1}^{I}\sigma^{i}y^{i}\in\mathcal{X}\text{ where }\sum\limits_{i=1}^{I}\sigma^{i}=1$
@@
Nikzad (2022) and Troyan (2024) consider a deterministic rank efficient assignment; that is, they focus on a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$. However, by Corollary 1, such $y^{\ast}$ must also be rank efficient in our definition; that is, $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$. Moreover, this result allows us to show that the mechanism introduced later is well-defined.
Finally, we introduce another technical result that is used to show our main results introduced later. Let
$\bar{k}\left(p_{a}\right)=\min\left\{k\in\left\{1,\cdots,\left|\mathcal{O}\right|\right\}\text{ }\left|\text{ }\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq k}q_{\bar{o}}\geq\left|\mathcal{A}\right|\right.\right\},$
which is well-defined because $\sum\nolimits_{\bar{o}\in\mathcal{O}}q_{\bar{o}}>\left|\mathcal{A}\right|$. This means that for a given $p_{a}$, there are sufficient number of copies of object types that ranked higher than or equal to $\bar{k}\left(p_{a}\right)$. Trivially, if $p_{a}^{k}=\varnothing$, then $\bar{k}\left(p_{a}\right)\leq k$. We have the following result.
@@
Let $x\in\mathcal{X}$ be an assignment that is not wasteful for $p$. Then, $x_{ao}=0$ for all $o$ such that $R(o,p_{a})>\bar{k}\left(p_{a}\right)$.
Then, if an assignment is not wasteful, then $a$ must be assigned to an object type that ranked higher than or equals to $\bar{k}\left(p_{a}\right)$. Typically, in that case, the object types that ranked lower than $\varnothing$ in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
@@ 3 Rank-minimizing Mechanism
Second, we introduce a fairness property. First, we introduce a well-known one. A mechanism $f$ is said to satisfy weak equal treatment of equals (hereafter weak ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}=p_{a^{\prime}}$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$. This requires that two agents revealing the same preference must have the same assignment in the ex-ante sense. This property is called just “equal treatment of equals” by many previous studies such as Bogomolnaia and Moulin (2001) and Troyan (2024). In this study, we mainly use a slightly stronger one. A mechanism $f$ is said to satisfy equal treatment of equals (hereafter ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $1,\cdots,\bar{k}\left(p_{a}\right)$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$ [^2] Han (2024) also considers a similar property, called symmetricity, which is slightly stronger than equal treatment of equals. Unlike in our model, they take into account the priorities of objects and define agents as equal only if they are equal with respect to the priorities of all objects.
By Lemma 2, as long as the assignment is not wasteful, any type that is ranked lower than $\bar{k}\left(p_{a}\right)$th in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than or equal to $\bar{k}\left(p_{a}\right)$th. Thus, if two individuals reveal rankings that are identical for positions at or above $\bar{k}\left(p_{a}\right)$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
@@
Let $\mathcal{Y}^{\ast}\left(p\right)\subseteq\mathcal{Y}$ be the set of deterministic rank efficient assignments; that is, $y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right),$ $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$. By Corollary 1, $\mathcal{Y}^{\ast}\left(p\right)\neq\emptyset$. Let
$f^{U}\left(p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}\text{ for all }p\in\mathcal{P}^{\left|\mathcal{A}\right|}$
@@
$RV\left(f^{U}\left(p\right),p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{1}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}RV\left(y^{\ast},p\right).$
Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank efficient for $p$. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the uniform rank-minimizing mechanism, this is one of the most common rank-minimizing mechanisms in the literature.
@@
Therefore, the uniform rank-minimizing mechanism is a fair rank-minimizing one.
We discuss strategic problems of fair rank-minimizing mechanisms. We call the strategy of revealing their true preference truth-telling. Featherstone (2020) shows that any rank-minimizing mechanism does not satisfy strategy-proofness. Therefore, in this paper, we consider a weaker strategic property.
We discuss whether truth-telling ($r_{a}$) can be strategically dominated by an untrue preference $p_{a}\neq r_{a}$. Let $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$. By Lemma 2, if $k>\bar{k}\left(p_{a}\right)$, then $p_{a}$ and $r_{a}$ are essentially equivalent under a fair rank-minimizing mechanism, because any result of a rank-minimizing mechanism is not wasteful. Thus, we focus only on the case where $k\leq\bar{k}\left(p_{a}\right)$ defined above; that is,
$\sum\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}\left(=\sum\limits_{\bar{o}:R(\bar{o},r_{a})<k}q_{\bar{o}}\right)<\left|\mathcal{A}\right|$
is satisfied for that $k$. In other word, we consider the case where the number of object types that are more preferable than the $k-1$th best one in $p_{a}$ is insufficient to distribute to all agents. In this case, for some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and some fair rank-minimizing mechanism, $a$ is assigned to an object type that is less preferable or equal to the $k$th best one.
@@
Suppose that $f$ is a fair rank-minimizing mechanism. Let $p_{a}\neq r_{a}$, and $k$ be the smallest integer such that $p_{a}^{k}\neq r_{a}^{k}$. If $k\leq\bar{k}\left(p_{a}\right)$, then there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that agent $a$ (with $r_{a}$) strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
By this lemma, we have the following result.
@@
This result implies that any fair rank-minimizing mechanism satisfies a weak strategic characteristic; that is, for each untrue preference $p_{a}$, either $\left(f\left(r_{a},p_{-a}\right)\right)_{ao}=\left(f\left(p_{a},p_{-a}\right)\right)_{ao}$ for all $o\in\mathcal{O}$ and all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ or there must exist $p_{-a}$ such that $a$ is worse off by revealing it instead of $r_{a}$. Thus, if a fair rank-minimizing mechanism is used, any (essential) strategic manipulation by an agent carries some level of risk. This characteristic is quite weak, but in the subsequent section, we show that this is not satisfied when the agents can refuse their assignment and instead obtain the null object type.
@@
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$, $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$ and $q_{o_{1}}=1$. Let $\psi=\left(o_{1},o_{2},\varnothing,o_{3}\right)$ and $\psi^{\prime}=\left(o_{1},o_{2},o_{3},\varnothing\right)$. A rank-minimizing mechanism $f$ is such that
@@
$f\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ $=$ $\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$
if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\psi^{\prime},\psi,\psi\right)$ nor $\left(\psi,\psi^{\prime},\psi\right)$ nor $\left(\psi,\psi,\psi^{\prime}\right)$. That is, if the best type and the second best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third best one obtains $o_{1}$ with probability 00. Trivially, $f$ satisfies weak ETE. However, whether $f$ satisfies ETE or not is dependent on $q_{o_{2}}$.
We assume that $r_{a_{1}}=\psi^{\prime}$ and this rank-minimizing mechanism is used. First, suppose $q_{o_{2}}=1$. In this case, $\bar{k}\left(\psi\right)=\bar{k}\left(\psi^{\prime}\right)=3$. Then, $r_{a_{1}}=\psi^{\prime}$ is not strategically dominated by $\psi$ and any other preference orders, because
@@
Second, suppose $q_{o_{2}}\geq 2$. In this case, $\bar{k}\left(\psi\right)=\bar{k}\left(\psi^{\prime}\right)=2$. Then, $a_{1}$ must have either $o_{1}$ or $o_{2}$ if it chooses either $\psi$ or $\psi^{\prime}$ and thus, the unique interest of $a_{1}$ is increasing the probability to obtain $o_{1}$. Thus, in this case, $r_{a_{1}}=\psi^{\prime}$ is strategically dominated by $\psi$. Therefore, even if a fair rank-minimizing mechanism is used, truth-telling may be strategically dominated by some other strategy. Moreover, in this case, $f$ does not satisfy ETE. First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing$,∅ , as long as the assignment is not wasteful. Thus, $\psi$ and $\psi^{\prime}$ are essentially the same in this case. However, $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}=0,$ but $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{2}}>0$ or $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{3}}>0$ must be satisfied. Therefore, in this case, $f$ does not satisfy ETE.